OpenStudy (anonymous):

Find the solution of y'' + 4y = 3t, y(0) = 0, y'(0) = 0

OpenStudy (abb0t):

Use variation of parameters or undefined coef method?

OpenStudy (anonymous):

\[y'' + 4y = 3t\]\[r^2 + 4 = 0\]\[r=\pm 2i\]\[0=\alpha, 2=\beta\]This is where I am

OpenStudy (anonymous):

I'm going to post my Yh if somebody could check it

OpenStudy (anonymous):

\[y_h=c_1e^{(0)t}\sin2t+c_2e^{(0)t}\cos 2t\]

OpenStudy (anonymous):

\[y_h=c_1\sin2t+c_2\cos 2t\]

OpenStudy (nubeer):

i dont think yh would be this.. i think sin and cos are in roots when complex roots.

OpenStudy (anonymous):

I have complex roots

OpenStudy (nubeer):

oops sorry ddnt saw it.

OpenStudy (abb0t):

Try using variation of parameters where Y1 = sin(2t) and Y2 = cos(2t). and use the wronskian. Then plug it into \[-Y _{1}\int\limits(\frac{ y _{2}g(t) }{ W })dt+Y_{2}\int\limits \frac{ y_{1}g(t) }{ W }dt\]

OpenStudy (abb0t):

where "W" is wronskian.

OpenStudy (anonymous):

honestly I haven't seen that. Can we stick with this method? I think my next step is to solve for c1 and c2 given my initial conditions

zepdrix (zepdrix):

No, before you can solve for c1 and c2, we need the particular solution.

OpenStudy (anonymous):

Oh ya, Yp

OpenStudy (abb0t):

Oh wait, zepdrix is right. I didn't see that it's complex. Sorry.

OpenStudy (nubeer):

try going with undetermined cofficient .. i think would be easy.

OpenStudy (anonymous):

I need to look through my notes a sec. I don't remember how to find Yp

OpenStudy (anonymous):

multiplicity (s) = 0 correct?

OpenStudy (abb0t):

yes. use undet coef and make a guess for 3t to solve and get ur particular solution

OpenStudy (abb0t):

Use Ct as your guess. Take y' and y'' and plug it into your solution and set it equal to Ct. Solve for C.

zepdrix (zepdrix):

Your particular solution will be of a form that matches the right side of your equation. In this case since we have a polynomial, our Yp will be a polynomial starting at degree 1,\[\huge y_p=At+B\]

OpenStudy (anonymous):

\[ay"+b'+cy=Ct^me^{rt}\]\[y_p=t^se^{rt}(At^m+...+A_0t^0)\]I got that too zep. \[y_p=At+B\]

OpenStudy (anonymous):

so my next step is to use Yp and the initial conditions to find A and B

zepdrix (zepdrix):

Mmmmm you should be able to find A and B without the initial conditions :D hmm

OpenStudy (anonymous):

\[y_p=At+B\]\[y_p'=A\]I'm stuck A=0 B=0 that doesn't seem right

zepdrix (zepdrix):

\[\large y_p=At+B, \quad y_p^{''}=0\]Plugging these into our equation gives us,\[\large y^{''}+4y=3t \quad \rightarrow \quad 0+4(At+B)=3t\]

OpenStudy (anonymous):

Is my Yp wrong because it is complex

zepdrix (zepdrix):

That is your Yc :) the complimentary solution. We don't want to mix that in with our Yp until we've found A and B.

zepdrix (zepdrix):

:o

OpenStudy (anonymous):

Ok I follow you @zepdrix

zepdrix (zepdrix):

\[\large y_c=c_1 \sin 2t+c_2 \cos 2t\]Yp will be of a form that matches the right hand side :D \[\large y_p=At+B\] Good now? k :3

zepdrix (zepdrix):

Just to remind you what we're trying to get :D\[\large y=y_c+y_p\]

OpenStudy (anonymous):

\[At+B=\frac{ 3 }{ 4 }t\]Now use my initial conditions

OpenStudy (anonymous):

B=0 A=3/4 ?

zepdrix (zepdrix):

Yah looks good! :)

zepdrix (zepdrix):

And NOWWWWW, after you get it into this form,\[\large y=y_c+y_p\]You can FINALLY use your initial conditions \:D/

OpenStudy (anonymous):

I want to go back to solving for A and B

OpenStudy (anonymous):

When I put Yp's into the original how do I solve for A and B. By equating terms?

zepdrix (zepdrix):

Yesss, remember how you do that in Partial Fractions? :) It's prretty much the same process.

OpenStudy (anonymous):

Ok that makes sense. I got lucky earlier :))

zepdrix (zepdrix):

Oh lol :D

OpenStudy (anonymous):

\[y=c_1 \sin 2t+c_2 \cos 2t+\frac{ 3 }{ 4 }t\]

zepdrix (zepdrix):

smooth.

OpenStudy (anonymous):

\[0=c_1 \sin 2(0)+c_2 \cos (0)t+\frac{ 3 }{ 4 }0\]\[0=c_2\]\[y'=2c_1 \cos2t+{3 \over 4}\]\[0=2c_1+{3 \over 4}\]\[c_1=-\frac{ 3 }{ 8 }\]\[y=-\frac{ 3 }{ 8 }\sin2t+\frac{ 3 }{ 4 }t\]^^Final Answer^^

zepdrix (zepdrix):

Yay good job \:D/

OpenStudy (anonymous):

Thanks man. I have my Diff EQ final Wed. When is yours? If I remember you're in this class right now too?

zepdrix (zepdrix):

Yah my Final is actually Wednesday also XD I have a pretty good teacher for this though, he keeps things rather simple :| it's almost been too easy at times. Spending most of my time preparing for Physics Final Exam tomorrow, and Calc 3 on Friday D: I'm not even worried about the DIff EQ Final.

OpenStudy (anonymous):

You shouldn't be. Strange how you're in Calc 3 at the same time as Diff EQ. My teacher is really smart but he just stares at the board and does problems. He does like 3 steps in his head at once sometimes and we all just look at eachother like WTF just happened. Good luck on your Physics final. I like physics